# Beam test results of a fully 3D-printed plastic scintillator particle detector prototype

### v2

@@ 4.3 Non-Uniformity in light response
We further investigated the light yield non-uniformity as a function of the ionization position within a single cube. As described earlier, the two hodoscopes with active area of 1.6 $\times$ 1.6 cm2 were placed at both sides of the prototype to cover the central row of cubes along the beam axis. A threshold at 300 ADC (about 4 p.e.) was applied to the hodoscope events to enhance the selection purity by rejecting the low light yield events, which are likely due to background signals. Tracks passing through all five cubes in the central row and both the hodoscopes were selected for analysis. By applying linear fits to the two hodoscope hits, tracks were reconstructed with higher spatial resolution. The middle point of the track segment in each cube was taken as the interacting point within the x-y plane perpendicular to the beam direction. The cross section of a single cube was split into 36 (6 $\times$ 6) bins, ensuring coverage for the high angle track segments. For each cube, the measured LY non-uniformity maps were generated separately for the two readout channels, perpendicular to the beam axis.
@@
Figure 8: The light yield non-uniformity map within a single cube readout by either the x (top) or y (bottom) fiber channels. The boundary of the scintillator cube is shown by the blue dashed line, while the position of the readout fiber is highlighted by the red dashed lines. The position of the other fiber is shown by the grey dashed lines.
Figure 8 presents the $6\times 6$-bin light yield distributions as a function of the track segment position within a single cube, for the x fibers (top) and the y fibers (bottom). Each map corresponds to the the same area as a center column of the prototype of size 1 cm2. The color of each bin represents the average light yield measured for tracks passing through that region in the cube. The edge of the cube is highlighted with blue dashed lines, while the active fiber for each map is highlighted in red, and the orthogonal fiber in gray. The maps clearly show higher light-yield regions around the WLS fibers. As expected, the closer the ionization position is to a fiber, the greater the detected light yield. Shadowing effects between the two WLS fibers are also visible: particles closer to one fiber produce less light detected by the orthogonal fiber. Defined as the relative deviation of the regional light yield from the overall average light yield, the light yield non-uniformity in the x channel ranges from -19% to 12%, while in the y channel it ranges from -16% to 12%. The overall root-mean-square (RMS) non-uniformity within a single cube was measured to be approximately 7%.
The analysis was extended to all five cubes in the central row, which were covered by the hodoscopes. For each cube, the RMS of the light yield non-uniformity was calculated from its non-uniformity map, and the results are summarized in Fig. 9. An average RMS non-uniformity of approximately 7% was observed, with low variation ($\sim$ 0.9%) across different cubes. This shows the good uniformity in the response of the 3D-printed scintillator cubes, as well as an excellent reproducibility.
@@
Figure 9: The light yield non-uniformity is derived as the RMS of the light yield measured at different particle interaction positions, as shown in Fig. 8. The measured non-uniformity of the five cubes in the central row covered by the hodoscoped is plotted above and the average is shown by the dashed red line (Top: x channel. Bottom: y channel). On average the non-uniformity within a single cube is about 7% with standard deviation of about 0.8% between different cubes.
@@ 5 Conclusions
The first prototype 3D-printed using the FIM technique, consisting of 125 plastic scintillator voxels, each of 1 $\times$ 1 $\times$ 1 cm3 size, and read out by two orthogonal WLS fibers, was successfully tested in a charged particle beam at the Proton-Synchrotron facility at CERN. In total 50 WLS fiber channels were instrumented, to provide 3D particle tracking. Two hodoscopes were installed at both side of the SuperCube, covering the whole cross section of the central cubes, enabling a high spatial resolution study of the light yield non-uniformity within a single voxel. A typical light yield of about 28 p.e. per channel was measured, being consistent with the result reported in [22] obtained with cosmic muons, and comparable with a sample manufactured by standard cast polymerization [23]. The cube-to-cube light leakage was measured to be 4% - 5% on average, meaning that in total about 24% - 30% of the detectable scintillating light escapes through reflective walls. The spatial non-uniformity of the detected light yield were measured from the center 5 voxels covered by the two hodoscopes. A light yield non-uniformity depending on the proximity of the interacting point to the WLS fibers is measured to be about 7% within a single cube. Less than 1% of variation has been observed by comparing the five cubes in the center column, showing the nice stability and reproducibility of the FIM manufactured plastic scintillator voxels. A new reflective filament is under developing and will potentially further reduce the light leakage of the prototype and increase the detected light yield.

### v3

@@ 4.4 Non-Uniformity in light response
We further investigated the LY non-uniformity as a function of the ionization position within a single cube. As described earlier, the two hodoscopes with active area of 1.6 $\times$ 1.6 cm2 were placed at both sides of the prototype to cover the central row of cubes along the beam axis. A threshold at 300 ADC (about 4 p.e.) was applied to the hodoscope events to enhance the selection purity by rejecting the low LY events, which are likely due to background signals. Tracks passing through all five cubes in the central row and both the hodoscopes were selected for analysis. By applying linear fits to the two hodoscope hits, tracks were reconstructed with higher spatial resolution. The middle point of the track segment in each cube was taken as the interacting point within the x-y plane perpendicular to the beam direction. The cross section of a single cube was split into 36 (6 $\times$ 6) bins, ensuring coverage for the high angle track segments. For each cube, the measured LY non-uniformity maps were generated separately for the two readout channels, perpendicular to the beam axis.
@@
Figure 8: The LY non-uniformity map within a single cube readout by either the x (top) or y (bottom) fiber channels. The boundary of the scintillator cube is shown by the blue dashed line, while the position of the readout fiber is highlighted by the red dashed lines. The position of the other fiber is shown by the grey dashed lines.
Figure 8 presents the $6\times 6$-bin LY distributions as a function of the track segment position within a single cube, for the x fibers (top) and the y fibers (bottom), averaged over all the five cubes of the center row. Each map corresponds to the the same area as a center column of the prototype of size 1 cm2. The color of each bin represents the average LY measured for tracks passing through that region in the cube. The edge of the cube is highlighted with blue dashed lines, while the active fiber for each map is highlighted in red, and the orthogonal fiber in gray. The maps clearly show higher light-yield regions around the WLS fibers. As expected, the closer the ionization position is to a fiber, the greater the detected LY. Shadowing effects between the two WLS fibers are also visible: particles closer to one fiber produce less light detected by the orthogonal fiber. Defined as the relative deviation of the regional LY from the overall average LY, the LY non-uniformity in the x channel ranges from -19% to 12%, while in the y channel it ranges from -16% to 12%. The overall root-mean-square (RMS) non-uniformity within a single cube was measured to be approximately 7%.
The analysis was extended to all five cubes in the central row, which were covered by the hodoscopes. For each cube, the RMS of the LY non-uniformity was calculated from its non-uniformity map, and the results are summarized in Fig. 9. An average RMS non-uniformity of approximately 7% was observed, with low variation ($\sim$ 0.9%) across different cubes. This shows the good uniformity in the response of the 3D-printed scintillator cubes, as well as an excellent reproducibility.
@@
Figure 9: The LY non-uniformity is derived as the RMS of the LY measured at different particle interaction positions, as shown in Fig. 8. The measured non-uniformity of the five cubes in the central row covered by the hodoscoped is plotted above and the average is shown by the dashed red line (Top: x channel. Bottom: y channel). On average the non-uniformity within a single cube is about 7% with standard deviation of about 0.8% between different cubes.
@@ 5 Conclusions
The first prototype 3D-printed using the FIM technique, consisting of 125 plastic scintillator voxels, each of 1 $\times$ 1 $\times$ 1 cm3 size, and read out by two orthogonal WLS fibers, was successfully tested in a charged particle beam at the Proton-Synchrotron facility at CERN. In total 50 WLS fiber channels were instrumented, to provide 3D particle tracking. Two hodoscopes were installed at both side of the SuperCube, covering the whole cross section of the central cubes, enabling a high spatial resolution study of the LY non-uniformity within a single voxel. A typical LY of about 28 p.e. per channel was measured, being consistent with the result reported in [22] obtained with cosmic muons, and comparable with a sample manufactured by standard cast polymerization [23]. The cube-to-cube light leakage was measured to be 4% - 5% on average, meaning that in total about 24% - 30% of the detectable scintillating light escapes through reflective walls. The spatial non-uniformity of the detected LY were measured from the center 5 voxels covered by the two hodoscopes. A LY non-uniformity depending on the proximity of the interacting point to the WLS fibers is measured to be about 7% within a single cube. Less than 1% of variation has been observed by comparing the five cubes in the center column, showing the nice stability and reproducibility of the FIM manufactured plastic scintillator voxels. A new reflective filament is under developing and will potentially further reduce the light leakage of the prototype and increase the detected LY. Work is in progress to add the WLS fiber holes for the third readout view.